# Thermoelectric Enhancement via Electronic and Phononic Channels in Staggered and Non-Staggered Dimerized Quantum Ring

### v1

@@ Abstract
Harnessing the quantum coherence and tunability of molecular-scale structures, we theoretically explore thermoelectric transport in ring-shaped molecular junctions featuring dimerized hopping integrals. By engineering alternating strong and weak bonds in both staggered and non-staggered configurations, we reveal a marked transmission asymmetry that drives a substantial enhancement in the thermoelectric figure of merit, ZTZT. To further steer transport behavior, we introduce controlled aperiodicity via site-energy modulations in unit cell format governed by the Aubry-André-Harper (AAH) potential, a quasiperiodic landscape that enables tunable localization-delocalization transitions. This interplay between hopping dimerization and AAH-type disorder gives rise to energy filtering effects and a rich spectrum where extended and critical states coexist, amplifying the Seebeck coefficient while preserving finite electrical conductance. Through a comprehensive non-equilibrium Green’s function analysis, we uncover how key device parameters, including disorder strength, dimerization amplitude, and lead-ring connectivity, collectively shape transport characteristics. Notably, asymmetric lead couplings are shown to enhance performance by leveraging quantum interference pathways. Our findings highlight a robust design strategy for optimizing nanoscale thermoelectric functionality, providing actionable insights for experimental realization in molecular electronic platforms.
@@ I Introduction
where $T$ is the absolute temperature, $S$ denotes the Seebeck coefficient, $G$ is the electrical conductance and $K_{\mathrm{el}}$ ($K_{\mathrm{ph}}$) represents the electronic (phononic) contribution to the total thermal conductance. Achieving a high ZTZT requires maximizing the power factor $S^{2}G$ while suppressing the parasitic heat flow carried by the phonons. However, in bulk conductors, the Wiedemann-Franz wf1 ; wf2 ; wf3 ; wf4 law enforces a proportionality between electronic thermal conductivity and electrical conductivity, making it difficult to enhance $G$ without increasing $K_{\mathrm{el}}$. Consequently, many state‐of‐the‐art bulk thermoelectric materials plateau at $ZT\sim 1$ near room temperature, and although some compounds exceed $ZT=2$ zt1 ; zt2 ; zt3 ; zt4 at elevated temperatures, scaling this performance to practical, large‐scale devices remains elusive.
@@
In this work, we analyze the thermoelectric properties of SSH‐type ring structures coupled to external leads in both symmetric and asymmetric configurations. Due to the ring topology, electrons can traverse multiple paths whose relative phases depend on contact positions. By attaching source and drain electrodes to inequivalent lattice sites, we engineer interferometric pathways that produce energy‐dependent transmission asymmetries, $\mathcal{T}(E)\neq\mathcal{T}(-E)$, necessary for obtaining a large Seebeck coefficient. We further superimpose the AAH potential [Eq. (2)] on each unit cell, allowing us to examine both staggered stag1 ; stag2 ; stag3 arrangements (where adjacent sites have opposite potential signs) and non‐staggered profiles (where the potential varies smoothly). By varying the modulation amplitude $\lambda$ and the dimerization ratio $t_{1}/t_{2}$, we systematically explore how band‐gap opening (due to hopping alternation) competes with correlated disorder (due to the AAH potential) to influence the coexistence of extended, critical, and localized eigenstates. Because ZTZT is highly sensitive to sharp features or gaps in the electronic density of states, this interplay can lead to pronounced enhancements in thermoelectric efficiency.
@@
Our study represents the first comprehensive investigation that simultaneously addresses (i) staggered versus non-staggered AAH potentials, (ii) dimerized hopping patterns characteristic of the SSH model, and (iii) geometric asymmetries in lead-ring couplings within a unified thermoelectric framework. We find that specific combinations of AAH modulation strength and dimerization ratio produce sharp transmission antiresonances antires and miniband structures, leading to peaks in the power factor $S^{2}G$ while suppressing total thermal conductance. These features reflect localization‐delocalization crossovers in the eigenstate spectrum, intimately related to the topological properties of the SSH Hamiltonian and the critical behavior of the AAH model. Our results provide clear design guidelines for molecular‐scale thermoelectric devices, showing that engineered bond alternation and correlated disorder in a ring geometry can achieve high ZTZT values. In particular, asymmetric ring‐lead connections further skew the transmission profile, enabling large Seebeck coefficients without significant loss of electrical conductance.
The remainder of this paper is organized as follows. In Sec. III we introduce the tight‐binding Hamiltonian for the dimerized SSH  ring with AAH modulations and outline the NEGF formalism for calculating charge and heat transport. Section IV presents a systematic analysis of transport characteristics as a function of dimerization ratio and AAH modulation strength comparing symmetric and asymmetric coupling geometries. We discuss the resulting electrical conductance, Seebeck coefficient, electronic thermal conductance, power factor, and figure of merit ZTZT, highlighting the roles of quantum interference and localization phenomena. Finally, Sec. V summarizes our main findings, elaborates on their implications for molecular‐scale thermoelectric applications, and outlines potential avenues for future research such as incorporating electron–phonon interactions and inelastic scattering effects to approach experimental realism.
@@ III.1 Quantum Ring Geometry and Tight-Binding Hamiltonian
We consider a one-dimensional quantum ring composed of NN unit cells arranged periodically to form a closed loop. Each unit cell contains two inequivalent atomic sites, labeled as $\alpha$ and $\beta$, as shown schematically in Fig. 2. The $\alpha$ and $\beta$ sites are depicted as distinct lattice points to reflect their differing local environments or on-site potentials. This bipartite lattice structure is a hallmark of the Su–Schrieffer–Heeger (SSH) model and introduces an intrinsic hopping asymmetry between sites, enabling the study of topological and localization phenomena in a unified platform.
@@ III.2 Lead Coupling and Thermoelectric Setup
To study thermoelectric transport, the ring is connected to two semi-infinite electrodes acting as thermal reservoirs, referred to as the source and drain maintained at temperatures $T\pm\Delta T/2$, respectively. The electrodes are modeled as one-dimensional tight-binding chains characterized by uniform site energies $\epsilon_{0}$ and nearest-neighbor hopping integrals $t_{0}$. These reservoirs are assumed to be reflectionless and ideal, ensuring unidirectional carrier injection and extraction. Coupling between the ring and the leads is restricted to specific lattice sites, typically the first and the NNth site of the ring to preserve quantum coherence and enforce boundary scattering conditions.
@@ III.3.2 Thermoelectric Response and Figure of Merit
Finally, the thermoelectric efficiency of the device is quantified by the dimensionless figure of merit ZTZT, defined as li1 ; li2
@@
A high value of ZTZT (typically $ZT\gtrsim 1$) indicates strong thermoelectric performance. In the quest for efficient nanoscale energy harvesters, achieving such values in quantum-coherent systems with tailored disorder or topological modulations is of significant contemporary interest.
@@ IV Numerical Results and Discussion
In this section, we present a comprehensive numerical investigation of the thermoelectric transport behavior of the proposed model. Our central focus is to identify the parameter regimes that favor enhanced thermoelectric performance, characterized by a high figure of merit (ZTZT). The study systematically explores how different physical parameters such as hopping amplitudes, onsite potential modulation, and structural configurations influence the electronic transport and thermoelectric response.
@@ IV.1 Energy-Dependent Transmission Function
These findings suggest that tuning the ratio $t_{1}/t_{2}$ in conjunction with the nature of the AAH potential (staggered vs non-staggered) provides a powerful knob for engineering transmission asymmetry an essential ingredient for achieving high Seebeck coefficients and, by extension, a large ZTZT. The subsequent sections will further analyze how these transmission features manifest in the thermoelectric performance metrics.
@@ IV.4 Electronic thermal transport characteristics as a function of Fermi energy
Figure 6 displays the variation of electronic thermal conductance $K_{\text{el}}$ as a function of the Fermi energy for different disorder and hopping configurations. In panel (a), corresponding to the non-staggered case with $t_{1}/t_{2}>1$, $K_{\text{el}}$ exhibits distinct behaviors for low and high disorder. While $W=0.3$ yields a broad distribution of $K_{\text{el}}$ near the band edges, $W=1.3$ results in mid-band thermal transport. Interestingly, these two curves exhibit limited overlap, effectively covering a wide energy range. The behavior generally parallels the trends in electronic conductance shown in Fig. 4(a), although deviations particularly for the high-disorder case highlight instances where the Wiedemann-Franz (WF) law does not hold. This breakdown is not uncommon in mesoscopic systems and is considered a key mechanism for enhancing the thermoelectric figure of merit ZTZT.
@@
IV.5 Variation of thermoelectric figure of merit ZTZT with Fermi energy
@@
Figure 7 presents the dependence of the thermoelectric figure of merit ZTZT on the Fermi energy $E_{F}$, under various hopping and disorder configurations in the Su-Schrieffer-Heeger (SSH) ring modified by a quasiperiodic Aubry-André-Harper (AAH) potential. The top row, panels (a) and (b), illustrates results for non-staggered onsite potentials, whereas the bottom row, panels (c) and (d), shows the staggered counterparts.
In the non-staggered case with $t_{1}/t_{2}>1$ [Fig. 7(a)], ZTZT reaches a remarkably high value of approximately 6 for the stronger disorder strength ($W=1.3$). This enhancement significantly surpasses the ZTZT values in the other configurations, where the maximum typically remains around 33 . Such a high ZTZT suggests an optimal regime where a combination of low electronic thermal conductance, moderate electrical conductance, and a sharp variation in the transmission function near the Fermi level leads to high Seebeck coefficient and efficient thermoelectric energy conversion.
@@
In comparison, the other three configurations non-staggered with $t_{1}/t_{2}<1$ [panel (b)] and both staggered cases [panels (c) and (d)]—display more modest ZTZT values, peaking around 3 for the same range of disorder. These results affirm that while the staggered arrangements still support finite thermoelectric efficiency, the synergy between the hopping configuration and the nature of the potential plays a decisive role in maximizing ZTZT. The non-staggered configuration with strong dimerization ($t_{1}/t_{2}>1$) emerges as the most favorable setting for thermoelectric optimization in these quasiperiodic nanostructures.
@@ IV.9 Disorder-induced modulation of phonon-mediated maximum thermal conductance
Interestingly, similar to earlier observations, the $t_{1}/t_{2}>1$ non-staggered case and the $t_{1}/t_{2}<1$ staggered case exhibit comparable behavior, again reflecting the transport symmetry present in this class of models. Suppression of $K_{\text{ph}}$ alongside preservation or enhancement of $S$ and $K_{\text{el}}$ is favorable for maximizing ZTZT, highlighting the utility of quasiperiodic SSH rings for phonon-glass electron-crystal-type thermoelectric engineering.
IV.10 Disorder-induced variation of thermoelectric efficiency (ZTZT)
@@
In the upper panel of Fig. 12, we examine the non-staggered case. In Fig. 12(a) For $t_{1}/t_{2}>1$, $ZT_{\text{max}}$ reaches values as high as 24. This remarkable enhancement can be physically attributed to a nontrivial interplay between disorder-induced localization and the resurgence of delocalized states at moderate disorder strengths. When the hopping ratio is inverted ($t_{1}/t_{2}<1$), [Fig. 12(b)], the system attains a peak ZTZT value of approximately 10, indicating substantial thermoelectric efficiency.
Such a nonmonotonic behavior in thermoelectric performance reflects the sensitive dependence of the transmission asymmetry on the disorder profile. In particular, the suppression of thermal conductivity due to localization, coupled with the retention or recovery of electrical conductance through resonant pathways, results in a highly favorable increase in ZTZT. This behavior is particularly evident in the $t_{1}/t_{2}>1$ non-staggered regime, where a reentrance of delocalized states enhances the asymmetry of the transmission function a key ingredient for high Seebeck response and thus elevated ZTZT.
In the lower panel of Fig. 12, we depict the $ZT_{\text{max}}$ variation for staggered configurations. For $t_{1}/t_{2}>1$ [Fig. 12(c)], the maximum ZTZT is found to be around 6, whereas for $t_{1}/t_{2}<1$ [Fig. 12(d)], it reaches nearly 9. These findings indicate that while the staggered system also exhibits reasonably good thermoelectric behavior, its efficiency is generally lower than its non-staggered counterpart, especially for the $t_{1}/t_{2}>1$ configuration.
@@ IV.11 Mapping thermoelectric performance across parameter space for non-staggered potentials
In Fig. 13(c), we present $K^{el}_{\text{max}}$, the electronic contribution to thermal conductance. Akin to $G_{\text{max}}$, $K^{el}_{\text{max}}$also displays reentrant characteristics. However, the qualitative behavior diverges in the $t_{1}/t_{2}>1$ regime, where $K^{el}_{\text{max}}$ remains relatively suppressed despite a rise in $G_{\text{max}}$. This deviation signals a violation of the Wiedemann–Franz (WF) law and hints at the potential for decoupled charge and heat transport an essential requirement for optimizing ZTZT.

### v2

@@ Abstract
Harnessing the quantum coherence and tunability of molecular-scale structures, we theoretically explore thermoelectric transport in ring-shaped molecular junctions featuring dimerized hopping integrals. By engineering alternating strong and weak bonds in both staggered and non-staggered configurations, we reveal a marked transmission asymmetry that drives a substantial enhancement in the thermoelectric figure of merit, $ZT$. To further steer transport behavior, we introduce controlled aperiodicity via site-energy modulations in unit cell format governed by the Aubry-André-Harper (AAH) potential, a quasiperiodic landscape that enables tunable localization-delocalization transitions. This interplay between hopping dimerization and AAH-type disorder gives rise to energy filtering effects and a rich spectrum where extended and critical states coexist, amplifying the Seebeck coefficient while preserving finite electrical conductance. Through a comprehensive non-equilibrium Green’s function analysis, we uncover how key device parameters, including disorder strength, dimerization amplitude, and lead-ring connectivity, collectively shape transport characteristics. Notably, asymmetric lead couplings are shown to enhance performance by leveraging quantum interference pathways. Our findings highlight a robust design strategy for optimizing nanoscale thermoelectric functionality, providing actionable insights for experimental realization in molecular electronic platforms.
@@ I Introduction
where $T$ is the absolute temperature, $S$ denotes the Seebeck coefficient, $G$ is the electrical conductance and $K_{\mathrm{el}}$ ($K_{\mathrm{ph}}$) represents the electronic (phononic) contribution to the total thermal conductance. Achieving a high $ZT$ requires maximizing the power factor $S^{2}G$ while suppressing the parasitic heat flow carried by the phonons. However, in bulk conductors, the Wiedemann-Franz wf1 ; wf2 ; wf3 ; wf4 law enforces a proportionality between electronic thermal conductivity and electrical conductivity, making it difficult to enhance $G$ without increasing $K_{\mathrm{el}}$. Consequently, many state‐of‐the‐art bulk thermoelectric materials plateau at $ZT\sim 1$ near room temperature, and although some compounds exceed $ZT=2$ zt1 ; zt2 ; zt3 ; zt4 at elevated temperatures, scaling this performance to practical, large‐scale devices remains elusive.
@@
In this work, we analyze the thermoelectric properties of SSH‐type ring structures coupled to external leads in both symmetric and asymmetric configurations. Due to the ring topology, electrons can traverse multiple paths whose relative phases depend on contact positions. By attaching source and drain electrodes to inequivalent lattice sites, we engineer interferometric pathways that produce energy‐dependent transmission asymmetries, $\mathcal{T}(E)\neq\mathcal{T}(-E)$, necessary for obtaining a large Seebeck coefficient. We further superimpose the AAH potential [Eq. (2)] on each unit cell, allowing us to examine both staggered stag1 ; stag2 ; stag3 arrangements (where adjacent sites have opposite potential signs) and non‐staggered profiles (where the potential varies smoothly). By varying the modulation amplitude $\lambda$ and the dimerization ratio $t_{1}/t_{2}$, we systematically explore how band‐gap opening (due to hopping alternation) competes with correlated disorder (due to the AAH potential) to influence the coexistence of extended, critical, and localized eigenstates. Because $ZT$ is highly sensitive to sharp features or gaps in the electronic density of states, this interplay can lead to pronounced enhancements in thermoelectric efficiency.
@@
Our study represents the first comprehensive investigation that simultaneously addresses (i) staggered versus non-staggered AAH potentials, (ii) dimerized hopping patterns characteristic of the SSH model, and (iii) geometric asymmetries in lead-ring couplings within a unified thermoelectric framework. We find that specific combinations of AAH modulation strength and dimerization ratio produce sharp transmission antiresonances antires and miniband structures, leading to peaks in the power factor $S^{2}G$ while suppressing total thermal conductance. These features reflect localization‐delocalization crossovers in the eigenstate spectrum, intimately related to the topological properties of the SSH Hamiltonian and the critical behavior of the AAH model. Our results provide clear design guidelines for molecular‐scale thermoelectric devices, showing that engineered bond alternation and correlated disorder in a ring geometry can achieve high $ZT$ values. In particular, asymmetric ring‐lead connections further skew the transmission profile, enabling large Seebeck coefficients without significant loss of electrical conductance.
The remainder of this paper is organized as follows. In Sec. III we introduce the tight‐binding Hamiltonian for the dimerized SSH  ring with AAH modulations and outline the NEGF formalism for calculating charge and heat transport. Section IV presents a systematic analysis of transport characteristics as a function of dimerization ratio and AAH modulation strength comparing symmetric and asymmetric coupling geometries. We discuss the resulting electrical conductance, Seebeck coefficient, electronic thermal conductance, power factor, and figure of merit $ZT$, highlighting the roles of quantum interference and localization phenomena. Finally, Sec. V summarizes our main findings, elaborates on their implications for molecular‐scale thermoelectric applications, and outlines potential avenues for future research such as incorporating electron–phonon interactions and inelastic scattering effects to approach experimental realism.
@@ III.1 Quantum Ring Geometry and Tight-Binding Hamiltonian
We consider a one-dimensional quantum ring composed of $N$ unit cells arranged periodically to form a closed loop. Each unit cell contains two inequivalent atomic sites, labeled as $\alpha$ and $\beta$, as shown schematically in Fig. 2. The $\alpha$ and $\beta$ sites are depicted as distinct lattice points to reflect their differing local environments or on-site potentials. This bipartite lattice structure is a hallmark of the Su–Schrieffer–Heeger (SSH) model and introduces an intrinsic hopping asymmetry between sites, enabling the study of topological and localization phenomena in a unified platform.
@@ III.2 Lead Coupling and Thermoelectric Setup
To study thermoelectric transport, the ring is connected to two semi-infinite electrodes acting as thermal reservoirs, referred to as the source and drain maintained at temperatures $T\pm\Delta T/2$, respectively. The electrodes are modeled as one-dimensional tight-binding chains characterized by uniform site energies $\epsilon_{0}$ and nearest-neighbor hopping integrals $t_{0}$. These reservoirs are assumed to be reflectionless and ideal, ensuring unidirectional carrier injection and extraction. Coupling between the ring and the leads is restricted to specific lattice sites, typically the first and the $N$th site of the ring to preserve quantum coherence and enforce boundary scattering conditions.
@@ III.3.2 Thermoelectric Response and Figure of Merit
Finally, the thermoelectric efficiency of the device is quantified by the dimensionless figure of merit $ZT$, defined as li1 ; li2
@@
A high value of $ZT$ (typically $ZT\gtrsim 1$) indicates strong thermoelectric performance. In the quest for efficient nanoscale energy harvesters, achieving such values in quantum-coherent systems with tailored disorder or topological modulations is of significant contemporary interest.
@@ IV Numerical Results and Discussion
In this section, we present a comprehensive numerical investigation of the thermoelectric transport behavior of the proposed model. Our central focus is to identify the parameter regimes that favor enhanced thermoelectric performance, characterized by a high figure of merit ($ZT$). The study systematically explores how different physical parameters such as hopping amplitudes, onsite potential modulation, and structural configurations influence the electronic transport and thermoelectric response.
@@ IV.1 Energy-Dependent Transmission Function
These findings suggest that tuning the ratio $t_{1}/t_{2}$ in conjunction with the nature of the AAH potential (staggered vs non-staggered) provides a powerful knob for engineering transmission asymmetry an essential ingredient for achieving high Seebeck coefficients and, by extension, a large $ZT$. The subsequent sections will further analyze how these transmission features manifest in the thermoelectric performance metrics.
@@ IV.4 Electronic thermal transport characteristics as a function of Fermi energy
Figure 6 displays the variation of electronic thermal conductance $K_{\text{el}}$ as a function of the Fermi energy for different disorder and hopping configurations. In panel (a), corresponding to the non-staggered case with $t_{1}/t_{2}>1$, $K_{\text{el}}$ exhibits distinct behaviors for low and high disorder. While $W=0.3$ yields a broad distribution of $K_{\text{el}}$ near the band edges, $W=1.3$ results in mid-band thermal transport. Interestingly, these two curves exhibit limited overlap, effectively covering a wide energy range. The behavior generally parallels the trends in electronic conductance shown in Fig. 4(a), although deviations particularly for the high-disorder case highlight instances where the Wiedemann-Franz (WF) law does not hold. This breakdown is not uncommon in mesoscopic systems and is considered a key mechanism for enhancing the thermoelectric figure of merit $ZT$.
@@
IV.5 Variation of thermoelectric figure of merit $ZT$ with Fermi energy
@@
Figure 7 presents the dependence of the thermoelectric figure of merit $ZT$ on the Fermi energy $E_{F}$, under various hopping and disorder configurations in the Su-Schrieffer-Heeger (SSH) ring modified by a quasiperiodic Aubry-André-Harper (AAH) potential. The top row, panels (a) and (b), illustrates results for non-staggered onsite potentials, whereas the bottom row, panels (c) and (d), shows the staggered counterparts.
In the non-staggered case with $t_{1}/t_{2}>1$ [Fig. 7(a)], $ZT$ reaches a remarkably high value of approximately 6 for the stronger disorder strength ($W=1.3$). This enhancement significantly surpasses the $ZT$ values in the other configurations, where the maximum typically remains around $3$ . Such a high $ZT$ suggests an optimal regime where a combination of low electronic thermal conductance, moderate electrical conductance, and a sharp variation in the transmission function near the Fermi level leads to high Seebeck coefficient and efficient thermoelectric energy conversion.
@@
In comparison, the other three configurations non-staggered with $t_{1}/t_{2}<1$ [panel (b)] and both staggered cases [panels (c) and (d)]—display more modest $ZT$ values, peaking around 3 for the same range of disorder. These results affirm that while the staggered arrangements still support finite thermoelectric efficiency, the synergy between the hopping configuration and the nature of the potential plays a decisive role in maximizing $ZT$. The non-staggered configuration with strong dimerization ($t_{1}/t_{2}>1$) emerges as the most favorable setting for thermoelectric optimization in these quasiperiodic nanostructures.
@@ IV.9 Disorder-induced modulation of phonon-mediated maximum thermal conductance
Interestingly, similar to earlier observations, the $t_{1}/t_{2}>1$ non-staggered case and the $t_{1}/t_{2}<1$ staggered case exhibit comparable behavior, again reflecting the transport symmetry present in this class of models. Suppression of $K_{\text{ph}}$ alongside preservation or enhancement of $S$ and $K_{\text{el}}$ is favorable for maximizing $ZT$, highlighting the utility of quasiperiodic SSH rings for phonon-glass electron-crystal-type thermoelectric engineering.
IV.10 Disorder-induced variation of thermoelectric efficiency ($ZT$)
@@
In the upper panel of Fig. 12, we examine the non-staggered case. In Fig. 12(a) For $t_{1}/t_{2}>1$, $ZT_{\text{max}}$ reaches values as high as 24. This remarkable enhancement can be physically attributed to a nontrivial interplay between disorder-induced localization and the resurgence of delocalized states at moderate disorder strengths. When the hopping ratio is inverted ($t_{1}/t_{2}<1$), [Fig. 12(b)], the system attains a peak $ZT$ value of approximately 10, indicating substantial thermoelectric efficiency.
Such a nonmonotonic behavior in thermoelectric performance reflects the sensitive dependence of the transmission asymmetry on the disorder profile. In particular, the suppression of thermal conductivity due to localization, coupled with the retention or recovery of electrical conductance through resonant pathways, results in a highly favorable increase in $ZT$. This behavior is particularly evident in the $t_{1}/t_{2}>1$ non-staggered regime, where a reentrance of delocalized states enhances the asymmetry of the transmission function a key ingredient for high Seebeck response and thus elevated $ZT$.
In the lower panel of Fig. 12, we depict the $ZT_{\text{max}}$ variation for staggered configurations. For $t_{1}/t_{2}>1$ [Fig. 12(c)], the maximum $ZT$ is found to be around 6, whereas for $t_{1}/t_{2}<1$ [Fig. 12(d)], it reaches nearly 9. These findings indicate that while the staggered system also exhibits reasonably good thermoelectric behavior, its efficiency is generally lower than its non-staggered counterpart, especially for the $t_{1}/t_{2}>1$ configuration.
@@ IV.11 Mapping thermoelectric performance across parameter space for non-staggered potentials
In Fig. 13(c), we present $K^{el}_{\text{max}}$, the electronic contribution to thermal conductance. Akin to $G_{\text{max}}$, $K^{el}_{\text{max}}$also displays reentrant characteristics. However, the qualitative behavior diverges in the $t_{1}/t_{2}>1$ regime, where $K^{el}_{\text{max}}$ remains relatively suppressed despite a rise in $G_{\text{max}}$. This deviation signals a violation of the Wiedemann–Franz (WF) law and hints at the potential for decoupled charge and heat transport an essential requirement for optimizing $ZT$.